OpenStudy (anonymous):

Con La regla del punto medio Encontrar: ∫_2^10▒√(x^3+1) dx,n=4

OpenStudy (anonymous):

∫_2^10 ▒ √(x^3+1) dx,n=4 [f(c1)+f(c2)+...+f(cn)]/n =(1/n)\[\sum_{1}^{n}\]∫_2^10▒√(x^3+1) \[\Delta x\]

OpenStudy (anonymous):

\[[f(c1)+f(c2)+...+f(cn)]/n =(1/n)\sum_{1}^{n} 2^ 10 \sqrt{x^3 +1} \Delta X\]